# Enhancing Memory and Imagination Consistency in Diffusion-based World Models via Linear-Time Sequence Modeling

### v1

@@ Appendix D Harmonizers Ablation Results
#Superhuman (↑) 0 N/A 9 11 11 13
Mean (↑) 0.000 1.000 1.124 1.364 1.674 1.866
@@
Appendix E Experimental Setup
We evaluate EDELINE on the Atari 100k benchmark (Schwarzer et al., 2023), which serves as the standard evaluation protocol in recent model-based RL literature for fair comparison. In addition, our experimental validation extends to ViZDoom (Kempka et al., 2016) and MiniGrid (Chevalier-Boisvert et al., 2023) environments to demonstrate broader applicability. To ensure statistical significance, all reported results represent averages across three independent runs. The Atari 100k benchmark (Schwarzer et al., 2023) encompasses 26 diverse Atari games that evaluate various aspects of agent capabilities. Each agent receives a strict limitation of 100k environment interactions for learning, in contrast to conventional Atari agents that typically require 50 million steps. EDELINE’s performance is evaluated against current state-of-the-art world model-based approaches, including DIAMOND (Alonso et al., 2024), STORM (Zhang et al., 2023), DreamerV3 (Hafner et al., 2024), IRIS (Micheli et al., 2023), TWM (Robine et al., 2023), and Drama (Wenlong et al., 2024). For evaluating 3D scene understanding capabilities, we employ VizDoom scenarios that demand sophisticated 3D spatial reasoning in first-person environments. This provides a crucial testing ground beyond the third-person perspective of Atari environments. Furthermore, the MiniGrid memory scenarios evaluate memorization capabilities through tasks that require information retention across extended time horizons.
Appendix F Theoretical Analysis of Mamba SSM Information Retention
F.1 Background and Motivation
Traditional deep reinforcement learning approaches rely on stacked frame inputs $s_{t}=[o_{t-k},\ldots,o_{t}]$ to capture temporal dependencies. This approach presents two fundamental limitations:
  1. 1.
    Fixed Memory Window: Information horizon is strictly bounded by window size $k$
  2. 2.
    Uniform Treatment: All observations within the window receive equal weighting regardless of importance
We propose replacing this architecture with Mamba State Space Models (SSM) utilizing hidden states $h_{t}$. Our analysis demonstrates superior information retention properties through rigorous mathematical proof.
F.2 Main Theorem and Implications
Theorem F.1 (Information Retention Superiority).
For any observation sequence $H_{t}=\{o_{1},\ldots,o_{t}\}$, with stacked frames $s_{t}$ of window size $k$ and Mamba SSM hidden state $h_{t}$, the following inequality holds:
$I(h_{t};H_{t})\geq I(s_{t};H_{t})+\Delta(t,k)$ (7)
where $\Delta(t,k)=t\epsilon-(k+1)H(o)$ becomes strictly positive when $t>\frac{(k+1)H(o)}{\epsilon}$.
This theorem establishes that Mamba’s information retention capability strictly surpasses traditional stacked frame approaches after a critical threshold, with the advantage growing linearly with sequence length. The term $\epsilon$ represents the minimum information retention rate per timestep, while $H(o)$ denotes the entropy of a single observation. The critical threshold $\frac{(k+1)H(o)}{\epsilon}$ represents the sequence length at which Mamba’s advantages begin to manifest.
F.3 Detailed Proof
The proof proceeds in three parts:
  1. 1.
    First, we establish an upper bound on information content in stacked frames
  2. 2.
    Then, we prove a lower bound on Mamba’s information retention
  3. 3.
    Finally, we combine these results to demonstrate the superiority condition
F.3.1 Part 1: Analysis of Stacked Frame Information Content
We begin by analyzing the information content of traditional stacked frames:
Definition F.2 (Stacked Frame Representation).
The stacked frame representation $s_{t}$ is defined as:
$s_{t}=[o_{t-k},\ldots,o_{t}]\in\mathcal{O}^{k+1}$ (8)
where $\mathcal{O}$ is the observation space.
Lemma F.3 (Stacked Frame Information Bound).
The mutual information between stacked frames and history is bounded by:
$I(s_{t};H_{t})\leq(k+1)H(o)$ (9)
Proof.
We apply the chain rule of entropy and use the properties of mutual information:
$I(s_{t};H_{t})$ $\leq H(s_{t})$ (Information bound) (10)
$=H(o_{t-k},\ldots,o_{t})$ (Definition of $s_{t}$) (11)
$\leq\sum_{i=0}^{k}H(o_{t-i})$ (Chain rule) (12)
$=(k+1)H(o)$ (Stationarity) (13)
The equality in the last step holds even the observations are dependent. This demonstrates the fundamental limitation of fixed-window approaches, as the information content is strictly bounded by the window size. ∎
Proposition F.4 (Information Loss Ratio).
The stacked frame architecture exhibits an information loss ratio of:
$L_{\text{stack}}=1-\frac{I(s_{t};H_{t})}{H(H_{t})}\geq 1-\frac{k+1}{t}$ (14)
Proof.
The loss ratio can be derived as follows:
$L_{\text{stack}}$ $=1-\frac{I(s_{t};H_{t})}{H(H_{t})}$ (Definition) (15)
$\geq 1-\frac{(k+1)H(o)}{tH(o)}$ (Using Lemma 1.3) (16)
$=1-\frac{k+1}{t}$ (Simplification) (17)
This reveals that the information loss grows linearly with the sequence length $t$, demonstrating the inefficiency of fixed-window approaches for long sequences. ∎
F.3.2 Part 2: Mamba SSM Information Retention Analysis
Definition F.5 (Mamba SSM State Update).
The Mamba SSM implements recursive state updates through:
$h_{t}=A_{t}h_{t-1}+B_{t}o_{t}$ (18)
where $h_{t}\in\mathbb{R}^{d}$ denotes the hidden state vector, $A_{t}\in\mathbb{R}^{d\times d}$ represents the state transition matrix, $B_{t}\in\mathbb{R}^{d\times m}$ is the input projection matrix, and $o_{t}\in\mathbb{R}^{m}$ is the observation vector.
Lemma F.6 (Information Evolution).
The mutual information between hidden state and history evolves according to:
$I(h_{t};H_{t})=I(h_{t};H_{t-1})+I(h_{t};o_{t}|H_{t-1})$ (19)
where $I(h_{t};o_{t}|H_{t-1})$ represents the new information gained from the current observation. The lemma aligns with Mamba’s recursive state update mechanism, which accumulates information adaptively.
Proof.
The proof follows from the chain rule of mutual information and the structure of the historical sequence:
$I(h_{t};H_{t})$ $=I(h_{t};\{H_{t-1},o_{t}\})$ (Definition of $H_{t}$)
$=I(h_{t};H_{t-1})+I(h_{t};o_{t}|H_{t-1})$ (Chain rule)
$=I(h_{t};H_{t-1})+\underbrace{[H(o_{t}|H_{t-1})-H(o_{t}|h_{t},H_{t-1})]}_{\text{Information gain from }o_{t}}$ (20)
The last term quantifies the reduction in uncertainty about $o_{t}$ provided by $h_{t}$, given the history $H_{t-1}$. This decomposition reveals how Mamba’s state update mechanism captures new information while preserving relevant historical context. ∎
Corollary F.7 (Information Monotonicity).
Under the Mamba update rule, the mutual information $I(h_{t};H_{t})$ is monotonically increasing in $t$:
$I(h_{t};H_{t})\geq I(h_{t-1};H_{t-1})+\epsilon_{t}$ (21)
where $\epsilon_{t}$ is the minimum information gain at time $t$.
Theorem F.8 (Information Growth).
Under Mamba’s selective memory mechanism:
$I(h_{t};H_{t})\geq\sum_{\tau=1}^{t}\epsilon_{\tau}\geq t\epsilon$ (22)
where $\epsilon=\min_{\tau}\epsilon_{\tau}>0$ is guaranteed by design.
Proof.
We proceed by induction on the sequence length $t$.
Base case $(t=1)$:
$I(h_{1};H_{1})=I(h_{1};o_{1})\geq\epsilon_{1}\geq\epsilon$ (23)
Inductive step: Assume the inductive hypothesis holds for $t-1$, i.e.,
$I(h_{t-1};H_{t-1})\geq\sum_{\tau=1}^{t-1}\epsilon_{\tau}$ (24)
Then for time step $t$:
$I(h_{t};H_{t})$ $=I(h_{t};H_{t-1})+I(h_{t};o_{t}|H_{t-1})$ (Chain rule) (25)
$\geq I(h_{t-1};H_{t-1})+\epsilon_{t}$ (DPI + Selectivity) (26)
$\geq\sum_{\tau=1}^{t-1}\epsilon_{\tau}+\epsilon_{t}=\sum_{\tau=1}^{t}\epsilon_{\tau}$ (Inductive hypothesis) (27)
$\geq t\epsilon$ (Definition of $\epsilon$) (28)
Therefore, by mathematical induction, the theorem holds for all $t\geq 1$. The linear growth in mutual information demonstrates Mamba’s ability to retain information over arbitrarily long sequences. ∎
F.3.3 Part 3: Comparative Analysis and Critical Threshold
Theorem F.9 (Information Advantage).
The information advantage $\Delta(t,k)$ of Mamba SSM over stacked frames is:
$\Delta(t,k)=I(h_{t};H_{t})-I(s_{t};H_{t})\geq t\epsilon-(k+1)H(o)$ (29)
This theorem provides practitioners with a quantitative decision tool: For any given window size $k$ and estimated retention rate $\epsilon$, one can compute the critical sequence length where Mamba becomes advantageous.
Proof.
Combining our previous results:
$\Delta(t,k)$ $=I(h_{t};H_{t})-I(s_{t};H_{t})$ (30)
$\geq t\epsilon-(k+1)H(o)$ (Theorems 2 and 3)
This advantage becomes positive when:
$t>\frac{(k+1)H(o)}{\epsilon}=t_{crit}$ (31)
F.4 Practical Implementation and Results
For practical implementation in Atari100k environments:
  • Standard window size: $k=4$
  • Empirically learned retention rate: $\epsilon\approx 0.2H(o)$
  • Resulting critical threshold: $t_{crit}=25$ frames
F.5 Optimality Conditions
The selection mechanism optimizes:
[…]
Subject to computational constraints:
$\text{rank}(A_{t})$ $\leq r$ (33)
$\|A_{t}\|_{F}$ $\leq\alpha$ (34)
where $r$ is the rank constraint and $\alpha$ bounds the Frobenius norm.
F.6 Key Theoretical Properties
  1. 1.
    Information Growth: Linear growth in mutual information with sequence length
  2. 2.
    Adaptive Memory: Input-dependent retention through learned matrices $A_{t},B_{t}$
  3. 3.
    Bounded Complexity: Linear computational scaling in sequence length
  4. 4.
    Provable Advantage: Quantifiable improvement over fixed-window approaches
These properties collectively address the core limitations of traditional approaches while maintaining computational tractability.
Appendix G Training time profile
Single update Time (ms) Detail (ms)
Total 548.6 148.6 + 400
   World model update 148.6 -
   Actor-Critic model update 400 15 $\times$ 24.4 + 34
    Imagination step (x 15) 24.4 17.2 + 6.5 + 0.7
     Next observation prediction 17.2 -
     Denoising step (x 3) 5.7 -
     Reward/Termination prediction 6.5 -
     Action prediction 0.7 -
    Loss computation and backward 34 -
Epoch Time (s) Detail (s)
Total 219.4 59.4 + 160
   World model 59.4 400 $\times$ 148.6 ×$\times 10^{-3}$
   Actor-Critic model 160 400 $\times$ 400 ×$\times 10^{-3}$
Run Time (days) Detail (days)
Total 2.9 2.5 + 0.4
   Training time 2.5 1000 $\times$ 219.4 / (24 $\times$ 3600)
   Other (collection, evaluation, checkpointing) 0.4 -
Module DIAMOND (ms/s/days) EDELINE (ms/s/days) Difference (%)
Single Update
Total 543 ms 548.6 ms +1.03%
   World Model Update 203 ms 148.6 ms -26.80%
   Actor-Critic Model Update 340 ms 400 ms +17.65%
Epoch
Total 217 s 219.4 s +1.11%
   World Model 81 s 59.4 s -26.67%
   Actor-Critic Model 136 s 160 s +17.65%
Appendix H edeline algorithm
We summarize the overall training procedure of edeline in Algorithm 1 below, which is modified from Algorithm 1 in Alonso et al. (2024). We denote as $\mathcal{D}$ the replay dataset where the agent stores data collected from the real environment, and other notations are introduced in previous sections or are self-explanatory.
Algorithm 1 EDELINE
1:  function training_loop():
2:     for epochs do
3:        collect_experience(steps_collect)
4:        for steps_world_model do
5:           update_world_model()
6:        end for
7:        for steps_actor_critic do
8:           update_actor_critic()
9:        end for
10:     end for
11:  end function
12:  function collect_experience($n$):
[…]
14:     for $t=0$ to $n-1$ do
15:        Sample $a_{t}\sim\pi_{\theta}(a_{t}\mid o_{t}^{0})$
16:        $o_{t+1}^{0},r_{t},d_{t}\leftarrow\texttt{env.step}(a_{t})$
17:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{o_{t}^{0},a_{t},r_{t},d_{t}\}$
18:        if $d_{t}=1$ then
[…]
20:        end if
21:     end for
22:  end function
23:  function update_world_model():
24:     Sample indexes :=$\mathcal{I}:=\{t,\;\dots,\;t+B+H-1\}$ // burn-in $B$ steps + imagination horizon $H$
25:     Sample sequence $(o_{i}^{0},\;a_{i},\;r_{i},\;d_{i})_{i\in\mathcal{I}}\sim\mathcal{D}$
26:     Initialize $h_{t-1}$ // MAMBA hidden state
27:     for $i\in\mathcal{I}$ do
28:        $h_{i}\leftarrow f_{\phi}(o_{i}^{0},\;a_{i},\;h_{i-1})$
29:        $\hat{r}_{i}\sim p_{\phi}(\hat{r}_{i}|h_{i})$
30:        $\hat{d}_{i}\sim p_{\phi}(\hat{d}_{i}|h_{i})$
31:     end for
32:     Compute reward modeling loss: $\mathcal{L}_{\mathrm{rew}}(\phi)\;=\;\sum_{i\in\mathcal{I}}\mathrm{CE}\bigl{(}\hat{r}_{i},\;r_{i}\bigr{)}$ // CE: cross-entropy loss
33:     Compute termination signal modeling loss: $\mathcal{L}_{\mathrm{end}}(\phi)\;=\;\sum_{i\in\mathcal{I}}\mathrm{CE}\bigl{(}\hat{d}_{i},\;d_{i}\bigr{)}$ // CE: cross-entropy loss
34:     Sample index $j$ $\sim$ $\text{Uniform}\{t+B,\dots,t+B+H-1\}$ // random sampling
35:     Sample $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ // log-normal sampling from EDM
36:     Define $\tau:=\sigma$ // identity schedule from EDM
37:     Sample $o_{j}^{\tau}\sim\mathcal{N}(o_{j}^{0},\;\sigma^{2}\mathbf{I})$ // add Gaussian noise
38:     Compute $\hat{o}_{j}^{0}=D_{\phi}\bigl{(}o_{j}^{\tau},\,\tau,\,o_{j-L}^{0},\dots,o_{j-1}^{0},h_{j-1}\bigr{)}$
39:     Compute observation modeling loss $\mathcal{L}_{\mathrm{obs}}(\phi)=\|\hat{o}_{j}^{0}-o_{j}^{0}\|^{2}$
40:     Update $\phi$ according to Eq. (4)
41:  end function
42:  function update_actor_critic():
43:     Sample initial buffer $(o_{t-B+1}^{0},\;a_{t-B+1},\;\dots,\;o_{t}^{0})\sim\mathcal{D}$
44:     // Burn in LSTM states $\pi_{\theta}$, $V_{\theta}$ and MAMBA states $f_{\phi}$ with the buffer
45:     for $i=t$ to $t+H-1$ do
46:        Sample $a_{i}\sim\pi_{\theta}(a_{i}\mid o_{i}^{0})$
47:        Compute $h_{i}\leftarrow f_{\phi}(o_{i},a_{i},h_{i-1})$
48:        Sample reward $r_{i}$, next observation $o_{i+1}^{0}$, and termination $d_{i}$ via $p_{\phi}$
49:     end for
50:     Compute $V_{\theta}(o_{i}^{0})$ for $i=t,\dots,t+H$
51:     Compute RL losses $\mathcal{L}_{V}(\theta)$ and $\mathcal{L}_{\pi}(\theta)$
52:     Update $\pi_{\theta}$ and $V_{\theta}$
53:  end function
Appendix I Hyperparameters
@@
Appendix J Actor-Critic Learning Objectives
We follow DIAMOND (Alonso et al., 2024) in the design of our agent behavior learning. Let $o_{t}$, $r_{t}$, and $d_{t}$ denote the observations, rewards, and boolean episode terminations predicted by our world model. We denote $H$ as the imagination horizon, $V_{\theta}$ as the value network, $\pi_{\theta}$ as the policy network, and $a_{t}$ as the actions taken by the policy within the world model.

### v2

@@ D.9.3 Effect of Harmonizers
#Superhuman ($\uparrow$) 0 N/A 9 11 11 13
Mean ($\uparrow$) 0.000 1.000 1.124 1.364 1.674 1.866
@@
D.9.4 Transformer-based Recurrent Embedding Module
To investigate the efficacy of Mamba compared to Transformer architecture, we implemented a Transformer-based variant of the Recurrent Embedding Module (REM). This ablation study aims to assess both performance and computational efficiency across memory-demanding environments, which provides critical insights into architecture selection for world modeling. We evaluated both architectures on the MiniGrid-MemoryS7, MiniGrid-MemoryS9, and Crafter environments, which specifically challenge a model’s ability to retain and utilize historical information.
Table 8: Performance and computational efficiency comparison between Transformer and Mamba-based REM architectures across memory-demanding environments. Values for MiniGrid environments represent success rates, while Crafter values show average return over three independent seeds after 1M environment steps. Mamba demonstrates comparable or superior performance with significantly better computational efficiency.
REM Architecture MiniGrid-MemoryS7 MiniGrid-MemoryS9 Crafter Training Time (per epoch)
Transformer 0.980 0.978 8.9 $\pm$ 0.6 293.0 sec
Mamba 0.981 0.982 11.5 $\pm$ 0.9 219.4 sec
As demonstrated in Table 8, both architectures achieve comparable performance on the simpler MiniGrid-Memory environments, with success rates that approach optimal performance. Nevertheless, MAMBA demonstrates a substantial $29.2\%$ improvement over Transformers in the more complex Crafter environment, which requires nuanced long-term memory and planning capabilities. Moreover, Mamba delivers approximately $25\%$ faster training times due to its linear time complexity compared to the quadratic complexity of Transformers.
These results substantiate our selection of Mamba as the foundation for EDELINE. Although both architectures perform comparably in simpler memory tasks, Mamba demonstrates superior performance in complex environments with intricate memory dependencies while maintaining superior computational efficiency. The combination of enhanced modeling capabilities and linear scaling with sequence length makes MAMBA particularly well-suited for world modeling applications that must process lengthy observation histories efficiently.
D.10 ViZDoom Environment Specifications
This appendix provides detailed specifications for the ViZDoom scenarios used in our experiments. We evaluate EDELINE on five key scenarios that test different agent capabilities in first-person 3D environments.
D.10.1 DeadlyCorridor
  • Reward Mechanism:
    • +1 for each enemy killed
    • +1 for armor acquisition
    • -1 for each damage instance received
D.10.2 HealthGathering
  • Objective: Survive by collecting medkits in toxic environment
  • Reward Mechanism:
    • +1 for each medkit collected
  • Evaluation Metric: Final health percentage (0-100%)
D.10.3 PredictPosition
  • Objective: Anticipate enemy movement to land delayed rocket hit
  • Reward Mechanism:
    • +1 for successful enemy kill
  • Evaluation Metric: Binary success if enemy eliminated
D.10.4 Basic
  • Objective: Eliminate stationary enemy with limited ammunition
  • Reward Mechanism:
    • +1 for successful kill
  • Evaluation Metric: Binary success if enemy eliminated
D.10.5 DefendCenter
  • Objective: Survive against infinite enemy waves
  • Reward Mechanism:
    • +1 per enemy killed
    • -1 for agent death
  • Evaluation Metric: Total enemies killed per episode
  • Observation space: (64,64,3)
  • Action space: Discrete movement and attack actions
D.11 Crafter Experiments Qualitative Results
Figure 13: Qualitative comparison of world model predictions in Crafter. We compare GT (Ground Truth), EDELINE, EDELINET (Transformer-based variant), and DIAMOND. The red boxes indicate prediction errors, the pink boxes show acceptable errors, and the green boxes highlight correct predictions of key elements.
In Fig. 13, we can observe comparative results between the different model architectures. In timesteps 1-10 (top row), all three methods demonstrate reasonable consistency in their predictions. Starting from timesteps 11-20 (middle row), notable differences emerge: DIAMOND fails to correctly represent crafted materials, and both DIAMOND and EDELINET incorrectly predict outcomes when the player chops trees and crafts new materials. EDELINE maintains higher prediction accuracy throughout these interactions. In timesteps 21-30 (bottom row), all methods generate acceptable predictions for previously unseen regions. While these differ from Ground Truth, they remain acceptable as they preserve gameplay-critical elements. During sleep actions, both EDELINE and DIAMOND fail to predict the sleeping state, which we classify as acceptable error given that sleep is a random event triggered independently of player actions. Most importantly, when the player returns to previously visited areas, only EDELINE successfully remembers and reproduces critical environmental features such as trees. EDELINET and DIAMOND both fail to recall the crafting table and trees outside the map border. This evidence demonstrates EDELINE’s superior memorization capability for maintaining environmental consistency over long horizons.
D.12 Hyperparameters
@@
Appendix E EDELINE Algorithm
We summarize the overall training procedure of EDELINE in Algorithm 1 below, which is modified from Algorithm 1 in [8]. We denote as $\mathcal{D}$ the replay dataset where the agent stores data collected from the real environment, and other notations are introduced in previous sections or are self-explanatory.
Procedure training_loop():
       for epochs do
             collect_experience(steps_collect)
             for steps_world_model do
                   update_world_model()
            for steps_actor_critic do
                   update_actor_critic()
Procedure collect_experience($n$):
[…]
       for $t=0$ to $n-1$ do
             Sample $a_{t}\sim\pi_{\theta}(a_{t}\mid o_{t}^{0})$
             $o_{t+1}^{0},r_{t},d_{t}\leftarrow\texttt{env.step(}a_{t}\texttt{)}$
             $\mathcal{D}\leftarrow\mathcal{D}\cup\{o_{t}^{0},a_{t},r_{t},d_{t}\}$
             if $d_{t}=1$ then
[…]
Procedure update_world_model():
       Sample indexes :=$\mathcal{I}:=\{t,\;\dots,\;t+T-1\}$ // sequence length $T$
       Sample sequence $(o_{i}^{0},\;a_{i},\;r_{i},\;d_{i})_{i\in\mathcal{I}}\sim\mathcal{D}$
       Initialize $h_{t-1}$ // MAMBA hidden state
       Parallel for $i\in\mathcal{I}$ do
             $h_{i}\leftarrow f_{\phi}(o_{i}^{0},\;a_{i},\;h_{i-1})$ // processed in parallel via MAMBA parallel scan
             $\hat{r}_{i}\sim p_{\phi}(\hat{r}_{i}|h_{i})$
             $\hat{d}_{i}\sim p_{\phi}(\hat{d}_{i}|h_{i})$
      Compute $\mathcal{L}_{\mathrm{rew}}(\phi)\;=\;\sum_{i\in\mathcal{I}}\mathrm{CE}\bigl{(}\hat{r}_{i},\;r_{i}\bigr{)}$ // CE: cross-entropy loss
       Compute $\mathcal{L}_{\mathrm{end}}(\phi)\;=\;\sum_{i\in\mathcal{I}}\mathrm{CE}\bigl{(}\hat{d}_{i},\;d_{i}\bigr{)}$ // CE: cross-entropy loss
       Sample index $j$ $\sim$ $\text{Uniform}\{t+B,\dots,t+T-1\}$ // burn-in $B$ steps
       Sample $\log(\sigma)\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ // log-normal sampling from EDM
       Define $\tau:=\sigma$ // identity schedule from EDM
       Sample $o_{j}^{\tau}\sim\mathcal{N}(o_{j}^{0},\;\sigma^{2}\mathbf{I})$ // add Gaussian noise
       Compute $\hat{o}_{j}^{0}=D_{\phi}\bigl{(}o_{j}^{\tau},\,\tau,\,o_{j-L}^{0},\dots,o_{j-1}^{0},h_{j-1}\bigr{)}$
       Compute observation modeling loss $\mathcal{L}_{\mathrm{obs}}(\phi)=\|\hat{o}_{j}^{0}-o_{j}^{0}\|^{2}$
       Update $\phi$ according to Eq. (5)
Procedure update_actor_critic():
       Sample initial buffer $(o_{t-B+1}^{0},\;a_{t-B+1},\;\dots,\;o_{t}^{0})\sim\mathcal{D}$
       // Burn in LSTM states $\pi_{\theta}$, $V_{\theta}$ and MAMBA states $f_{\phi}$ with the buffer
       for $i=t$ to $t+H-1$ do
             Sample $a_{i}\sim\pi_{\theta}(a_{i}\mid o_{i}^{0})$
             Compute $h_{i}\leftarrow f_{\phi}(o_{i},a_{i},h_{i-1})$
             Sample reward $r_{i}$, next observation $o_{i+1}^{0}$, and termination $d_{i}$ via $p_{\phi}$
      Compute $V_{\theta}(o_{i}^{0})$ for $i=t,\dots,t+H$
       Compute RL losses $\mathcal{L}_{V}(\theta)$ and $\mathcal{L}_{\pi}(\theta)$
       Update $\pi_{\theta}$ and $V_{\theta}$
Algorithm 1 EDELINE
Appendix F Actor-Critic Learning Objectives
We follow DIAMOND [8] in the design of our agent behavior learning. Let $o_{t}$, $r_{t}$, and $d_{t}$ denote the observations, rewards, and boolean episode terminations predicted by our world model. We denote $H$ as the imagination horizon, $V_{\theta}$ as the value network, $\pi_{\theta}$ as the policy network, and $a_{t}$ as the actions taken by the policy within the world model.